b. If one car is randomly chosen, find the probability that it is traveling more than 78 mph


1
Expert's answer
2022-03-01T06:04:06-0500

Let "X" be the speed for a randomly selected car: "X\sim N(\mu, \sigma^2)."

Given "\mu=79\ mph, \sigma=10\ mph."


"P(X>78)=1-P(X\le78)"

"=1-P(Z\le \dfrac{78-79}{10})=1-P(Z\le-0.1)"

"\approx0.5398"